• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi/Miha

    Log entry time 18:51:31 on June 10, 2009

    Entry number 280444

    This entry is a followup to: 272648

    Followups:

    keyword=Replaced damaged jumper fiber for the comet lasers

    During the time I was checking the fiber temperature after we switched to
    the longitudinal pumping, I found out the connections between the jumper
    fibers from Rutgers and Jlab lasers to the 75 long fibers were
    very hot. After talked to RC, we decided to take a look by turning off
    these lasers.

    It turned out to be that the jumper of the Rutgers and Jlab laser were
    badly damaged. We decided to change black jumper fiber cables (that lead
    from the laser diode to the metal cables that are going to the Hall) for
    these two lasers.

    When we were attaching cables (from laser controller
    and temperature controller) back to diode enclosure one of the screws on
    the connector (on the cable from laser controller) broke. We decided to
    replace this cable with the cable from the new, recently bought laser.

    After that we also did the laser power test, to make sure, that lasers
    are working properly. In the end we connected new fibers to the fibers
    leading to the Hall and turned on the lasers.